how many conversation will there be at a party of 50 people. (each person only talk to each other once.)
Found 2 solutions by venugopalramana, cleomenius:
Answer by venugopalramana(3286)

THE ANSWER IS 50C2..HOPE YOU KNOW COMBINATIONS.
=50*49/2=1225

Answer by cleomenius(959)

The result should work out as 49 + 48 + 47 + ,,, +3 + 2 + 1
There should be a total of 24 pairs of 50, with each pair being 50.
Plus, there should be a single value of 25 left over ( the middle value).
Therefore the result should be (24)(50) + 25 = 1225